OpenStudy (lucidiousxiv):

-=[Medal]=- What are the foci of the graph x2 – 4y2 = 16? (+/- 4.5, 0) (+/- 3.9, 0) (+/- 4.1, 0) (+/- 2.2, 0)

OpenStudy (lucidiousxiv):

@iGreen

OpenStudy (anonymous):

its either a or c? what did you get

OpenStudy (anonymous):

i think its c

OpenStudy (anonymous):

not too sure

OpenStudy (ybarrap):

The equation of a hyperbola is $$ {x^2\over a^2}-{y^2\over b^2}=1 $$ In your equation a=4 and b = 2 Also, the distance \(c\) from the center to either focus is $$ c^2=a^2+b^2 $$  Just plug in a and b to get c, the x-coordinate of your focus. Since this is an east-west parabola symmetric about the x-axis, the y-coordinate is zero. see - http://en.wikipedia.org/wiki/Hyperbola http://www.wolframalpha.com/input/?i=x%5E2-4y%5E2+%3D+16